OpenStudy (anonymous):

Jeff needs to buy food and medication for six cats and one dog. The cost of food and medication for a cat is $5 less than two-thirds of the cost of food and medication for a dog. If Jeff spends a total of $195, the amount he spent on food and medication for the dog is $

OpenStudy (aripotta):

cost for dog = d cost for cat = 2/3d - 5 there are six cats so 6(2/3d - 5) he spent a total of $195 so the equation will be equal to $195 so 6(2/3d - 5) + d = 195 can you solve for d now?

OpenStudy (aripotta):

@zxcvbnm08765 you there?

OpenStudy (anonymous):

im still lost @AriPotta

OpenStudy (aripotta):

i'm gonna walk my dog real quick. i'll be back in about 10 minutes to help you

OpenStudy (anonymous):

k @AriPotta

OpenStudy (aripotta):

alrighty. what are you confused with?

OpenStudy (aripotta):

let me break down the question for you. we're going to assign the variable "d" for the price of the dog's food and medications. the question says "The cost of food and medication for a cat is $5 less than two-thirds of the cost of food and medication for a dog" so 2/3d means 2/3 the cost of dog food and meds. but it also says it's $5 less than the cost of the dog stuff so we have to subtract 5 from 2/3d so... cost of dog = d cost of cat = 2/3d - 5 cost of 6 cats = 6(2/3d - 5) the question says "Jeff spends a total of $195" so the price of both the dog and cat stuff is going to be equal to 195 so... 6(2/3d - 5) + d = 195 from there, it's simple algebra to solve for d. can you do that?

OpenStudy (anonymous):

its 45 @AriPotta

OpenStudy (aripotta):

correct :)